# E4S: Fine-grained Face Swapping via Editing With Regional GAN Inversion

### v2

@@ A.0.1 Loss for the RGI
A.0.2 Loss for the Face Re-coloring Network $B_{\psi}$
Reconstruction Loss. During training, given $I^{\prime}_{\rm{A}}$ and and $I^{\prime}_{\rm{AG}}$, the re-coloring network learns to predict the re-colored $\tilde{I}^{*}$. The predicted $\tilde{I}^{*}$ should be fully consistent with the $I_{\rm{A}}$ (which equals to the flipped $I_{\rm{A}}^{\prime}$). To this end, we use a reconstruction loss consisting of an L2 loss and an LPIPS loss as the training objective:
$\mathcal{L_{\psi,\rm{rec}}}=\lambda_{\psi,1}\|I_{\rm{A}}-\tilde{I}^{*}\|^{2}_{2}+\lambda_{\psi,2}\|\rm{V}(I_{\rm{A}})-\rm{V}(\tilde{I}^{*})\|^{2}_{2},$ (25)
@@ A.0.2 Loss for the Face Re-coloring Network $B_{\psi}$
A.0.3 Details of Training the Face Inpainting Network $P_{\tau}$
Training Settings and Loss Function. Similar to training re-coloring network $B_{\psi}$, training inpainting network $P_{\tau}$ requires paired data which is hard to collect. Therefore, we train $P_{\tau}$ in a self-supervised scheme. Given a face $I$, we generate random mismatch masks and edit its skin belonging to these mismatch regions based on the editing ability of our RGI method, obtaining an edited face $I_{\rm{edit}}$. Then we take $I_{\rm{edit}}$ as the training input and $I$ as the ground-truth supervision with a reconstruction loss:
$\mathcal{L_{\tau,\rm{rec}}}=\lambda_{\tau,1}\|I_{\rm{A}}-\tilde{I}^{*}\|^{2}_{2}+\lambda_{\tau,2}\|\rm{V}(I_{\rm{A}})-\rm{V}(\tilde{I}^{*})\|^{2}_{2},$ (26)
@@ Appendix B Video Face Swapping Results
The proposed E4S framework can be used to conduct video face swapping. Specifically, following STIT [63], we first crop and align the source image and the target video in advance, thus resulting in the source face S and an n-frame target face video $\{T_{i}\}_{i=1}^{n}$. Then, as described in Sec. 3.1 in our main paper, we can perform face re-enactment on the source S and make it have a similar pose and expression as each target frame. In this way, we can obtain a sequence of driven video $\{D_{i}\}_{i=1}^{n}$. Next, we can perform face swapping between each driven and target pair $\{(D_{i},T_{i})\}_{i=1}^{n}$.
@@
We compare our results with FaceShifter [9] and HiRes [3] in Fig. 13, where the whole video is on our project page. HiRes struggles to generate a wink in the swapped video. Besides, FaceShifter sometimes fails to transfer the source identity. As contrast, our results show better visual quality and temporal consistency.
@@ Appendix C User-interface System For Image and Video Face Swapping
We develop a user-interface system to perform image and video face swapping. A screenshot is shown in Fig. 14. Given a source image and a target image or video, our user-interface system can produce a high-quality swapped image or video. An interactive face swapping demo can also be found on our project page.

### v3

@@ A-1 Loss for the RGI
A-2 Loss for the Face Re-coloring Network $B_{\psi}$
Reconstruction Loss. During training, given $I^{\prime}_{\rm{A}}$ and and $I^{\prime}_{\rm{AG}}$, the re-coloring network learns to predict the re-colored $I_{\rm{rec}}$. The predicted $I_{\rm{rec}}$ should be fully consistent with the $I_{\rm{A}}$ (which equals to the flipped $I_{\rm{A}}^{\prime}$). To this end, we use a reconstruction loss consisting of an L2 loss and an LPIPS loss as the training objective:
$\mathcal{L_{\psi,\rm{rec}}}=\lambda_{\psi,1}\|I_{\rm{A}}-I_{\rm{rec}}\|^{2}_{2}+\lambda_{\psi,2}\|{\rm{V}}(I_{\rm{A}})-{\rm{V}}(I_{\rm{rec}})\|^{2}_{2},$ (23)
@@ A-2 Loss for the Face Re-coloring Network $B_{\psi}$
A-3 Details of Training the Face Inpainting Network $P_{\tau}$
Training Settings and Loss Function. Similar to training re-coloring network $B_{\psi}$, training inpainting network $P_{\tau}$ requires paired data which is hard to collect. Therefore, we train $P_{\tau}$ in a self-supervised scheme. Given a face $I$, we generate random mismatch masks around the face contour and edit these regions based on the editing ability of our RGI method. The edited face $I_{\rm{edit}}$ have an inconsistent face shape with the original one. Then we adopt $I_{\rm{edit}}$ as the training input and $I$ as the ground-truth supervision with a reconstruction loss, enforcing the inpainting network to predict a shape-consistent face result $I_{\rm{inp}}$ under the guidance of the aforementioned random mismatch mask:
$\mathcal{L_{\tau,\rm{inp}}}=\lambda_{\tau,1}\|I-I_{\rm{inp}}\|^{2}_{2}+\lambda_{\tau,2}\|{\rm{V}}(I)-{\rm{V}}(I_{\rm{inp}})\|^{2}_{2},$ (24)
@@ A-3 Details of Training the Face Inpainting Network $P_{\tau}$
Ap-Fig. 14: Video face swapping comparisons of our results with FaceShifter [5] and HiRes [45]. Our method shows the better capability of source identity transferring and target attribute preservation (e.g., pose, expression, wink, lighting). Their visual quality and temporal consistency also inferior to our method.
Ap-Fig. 15: A screenshot of our interactive image and video face swapping system. The button in the upper left corner can choose whether to swap face in images or videos.
@@ Appendix B Video Face Swapping Results
The proposed E4S framework can be used to conduct video face swapping. Specifically, following STIT [67], we first crop and align the source image and the target video in advance, thus resulting in the source face S and an n-frame target face video $\{T_{i}\}_{i=1}^{n}$. Then, as described in Sec. III-A in our main paper, we can perform face re-enactment on the source S and make it have a similar pose and expression as each target frame. In this way, we can obtain a sequence of driven video $\{D_{i}\}_{i=1}^{n}$. Next, we can perform face swapping between each driven and target pair $\{(D_{i},T_{i})\}_{i=1}^{n}$.
@@
We compare our results with FaceShifter [5] and HiRes [45] in Ap-Fig. 14, where the whole video is on our project page. HiRes struggles to generate a wink in the swapped video. Besides, FaceShifter sometimes fails to transfer the source identity. As contrast, our results show better visual quality and temporal consistency.
@@ Appendix C User-interface System For Image and Video Face Swapping
We develop a user-interface system to perform image and video face swapping. A screenshot is shown in Ap-Fig. 15. Given a source image and a target image or video, our user-interface system can produce a high-quality swapped image or video. An interactive face swapping demo can also be found on our project page.